OpenStudy (anonymous):

What is the inverse of the given relation? y = 3x + 12

OpenStudy (anonymous):

Will give medals, please help!

OpenStudy (jokersmart):

Simplifying Y = 3x + 12 Reorder the terms: Y = 12 + 3x Solving Y = 12 + 3x Solving for variable 'Y'. Move all terms containing Y to the left, all other terms to the right. Simplifying Y = 12 + 3x

OpenStudy (anonymous):

switch x and y so x = 3y +12 solve for y 3y=-12

OpenStudy (anonymous):

y^-1 = -12/3 = -4

OpenStudy (anonymous):

Thank you everyone!

OpenStudy (anonymous):

yw

OpenStudy (adi3):

did u get it